Work Order Lookup
| Date | WO # | Location | Description | Requester | Status |
|---|---|---|---|---|---|
| 8/29/2022 | 218465 | 3600 HOWE | Attn: Jeremy Forristall -exhaust fan EX25 has new VFD could not get it to start in auto had to switch it local to get it started and switch it back to auto but was not able to get it to take off from stopped in auto | Jeffrey Lamb | Assigned |
| 1/18/2023 | 252666 | GENERAL HOWE | VFD 5 for EF2 needs replaced. Attention Jeremy F | Dustin Limoges | Materials-Awaiting |
| 3/2/2023 | 265265 | 1350 HOWE | Clean black mold from Howe Hall entryway. | James Wellman | Assigned |
| 4/3/2023 | 272160 | 0110 HOWE | There is a on-going issue in the the 1140-4 hallway of water leaking. The hallway gets wet and fans are run to dry it up. The small never goes away. This happens anytime there is signicant rain. Friday's storm was not exception. Attached is a photo of the likely cause. I would like to request the cause be addressed. | Arlene Grebasch | Open |
| 12/13/2023 | 329885 | 3200 HOWE | Replace vfd 12 on air handler 2 | Jeremy Forristall | Assigned |
| 1/25/2024 | 337139 | 0612 HOWE | Howe mech room 0612. The steam water heater (WH1) is leaking at the bottom appears to be the steam return line-piping is very rusty. | Jeffrey Ellsworth | Assigned |
| 1/26/2024 | 337315 | HOWE | Please deliver bleachers from Lied Rec Services to Howe Hall Atrium and place them on the floor between pillars and opposite side of facing bleachers; for Lego League., and pick up Monday, January 29, 2024. EVENT: Lego League, Friday, January 26 - 29, 2024, Howe Hall Atrium, CONTACT: Sam Franzen, sfranzen@iastate.edu; 515-296-3434 FUNDING: PG101480 DD17467 University event | Barbara Schumacher | Open |
| 4/30/2024 | 357921 | GENERAL HOWE | Air handler 3 the motor will run and the it trips out at the MCC. Motor needs checked and also the amp draw. Attention Jeremy F | Dustin Limoges | Materials-Awaiting |
| 5/24/2024 | 364056 | GENERAL HOWE | Storm damage - small trees - maybe saveable? | Arlene Grebasch | Assigned |
| 8/19/2024 | 383518 | 2624F HOWE | The wire management system on the bottom of the table has lost its adhesion (it is now on the ground) and needs to be re-adhered to help with the cable management, and there is less chance of people tripping and messing up the system. Additional location information: 2624F Howe Hall - (BIG) Conference Room - CELT Suite Work tags to attach: PG113293 | Laura Bestler | Assigned |
| 8/20/2024 | 383771 | 1010 HOWE | The fascia behind the Department of Aerospace Engineering is damaged and needs to be repaired which then means it will need to be repainted. Damage is obvious when you are in front of area. | Arlene Grebasch | Assigned |
| 1/24/2025 | 416680 | 3600 HOWE | EF-21 will not work, don't try and run it without Joe, Josh or mike being available to open dampers, ATTN Jermey and Dana | Joe Flaherty | Assigned |
| 2/28/2025 | 424080 | 3200 HOWE | Scope of work repair or replace Howe AHU 4 VFD | Jeremy Forristall | Assigned |
| 4/7/2025 | 434447 | 3200 HOWE | Replace VFD on AHU-2. RF-2B | Jeremy Forristall | Assigned |
| 5/8/2025 | 441644 | 1280 HOWE | The knee wall has gotten very dirty and damaged and needs to be repaired and painted including the white metal railings. If possible, it would be great to put a protective panel on the corners (directly out from the overhead and next to the over head door) since this area gets a lot of bumps due to the loading and unloading of freight in the area. Pictures of current condition attached. | Arlene Grebasch | Assigned |
| 5/19/2025 | 442403 | HOWE | There is a “moon” tree being planted by Howe. The President’s office is paying for the install, the concrete edging, and the fencing to protect the tree. | Rhonda Martin | Assigned |
| 5/19/2025 | 442537 | HOWE | Elevator is stuck on the first floor and unresponsive. | James Benson | Assigned |
| 6/5/2025 | 447629 | 1630 HOWE | Repair Fire Alarm Inspection deficiencies: "Fire door by 0304 is missing it’s closer Fire door west of 0224 is missing its closer Smoke detector: Room 1304, Located above ceiling in room 1373B-1, 2ND LVL AREA ""C"" ROOM 2624H,K,C,G, Needs remounted, located above ceiling LWR LVL ""B"" VEST 1140D - SMOKE DETECTOR: Inside garbage bag above ceiling" ATTN: Tim Kelley | Sailesh Gautam | Assigned |
| 6/12/2025 | 448581 | 1246 HOWE | 1246 room inspection noted the following: - touch-up spackle/paint under instructor marker wall, on lower wall half. - remove black marker graffiti on classroom (back) fabric wall. - sconce/globe wall needs wall fabric repair in a couple of spots. | Dale Headrick | Assigned |
| 6/19/2025 | 449255 | 1328E HOWE | We are taking over this area from IMSE and will be using this room as an ITAR secure room. We need a University monitored lock on the door and are interested in moving the lock from the exterior hall door to 1328-3. However, we don't know if this is possible or advisable. It might be better to install new. Before action is taken we'd like to discuss the possibilities and costs of this action with the appropriate person. We need to wait until 7/1/25 for the work to be done. Discussion can happen at the earliest convenient time. | Arlene Grebasch | Open |
| 6/24/2025 | 450879 | 2271 HOWE | Randy Tabber borrowed my upright to cover Andy's area, Brock and returned it with clogged hair in the agitator. I removed the hair, but the belt on the agitator is still running rough. | Nickolas Thielen | Materials-Awaiting |
| 7/7/2025 | 457820 | 0612 HOWE | Hire Delta Industries to perform major PMs on Air Compressor Unit ACU2. | Sailesh Gautam | Open |
| 7/7/2025 | 457828 | 0612 HOWE | Hire Delta Industries to perform major PMs on Air Compressor Unit ACU3. | Sailesh Gautam | Open |
| 7/7/2025 | 457831 | 0612 HOWE | Hire Delta Industries to perform major PMs on Vacuum Pump. | Sailesh Gautam | Open |
| 7/10/2025 | 458655 | 2271 HOWE | This carpet wand pictured in room 2271 doesn't squirt fluid for carpet extractor when hooked up to the Karcher extractor. If Brock gets it to work, then it needs blue hoses to match it as that has been lost since I have started here at the University. | Nickolas Thielen | Materials-Awaiting |
| 7/10/2025 | 458722 | 1214 HOWE | Brock we need a wand for the pictured carpet extractor. One that could cover a swath like the pictured wand that won't fit the smaller extractor also pictured. | Nickolas Thielen | Materials-Awaiting |
| 7/11/2025 | 458853 | HOWE | The alarms are going off - the panel is reading "Communication Failure w/ master" | Arlene Grebasch | Open |
| 7/14/2025 | 459087 | HOWE | Please clean out the sand under the walking grates at the front entrance. | Christopher Young | Open |
| 7/16/2025 | 459515 | 1635 HOWE | Autoscrubber circle brush is off. Looks like the pin is missing to hold the brush in place. Could you please replace the brush and pin Thanks in advance Brock | Christi Johnson-Bascomb | Materials-Awaiting |
| 7/16/2025 | 459575 | 1214 HOWE | Light super sucker coworker states it doesn't suck up. The darker super sucker needs new rubber squeegee pieces there torn. Thanks in advance for fixing | Christi Johnson-Bascomb | Materials-Awaiting |
| 7/16/2025 | 459577 | 1214 HOWE | The long handle to the carpet fan is needing tighten or replaced. | Christi Johnson-Bascomb | Materials-Awaiting |
| 7/17/2025 | 459689 | HOWE | Repair Wet Sprinkler deficiencies: One fire sprinkler head above the virtual projector at the top of the atrium area is taped and one is painted. The ceiling is too high to determine the type of sprinkler head needing to be replaced for the painted one and that area looks very difficult to access. A site survey is needed to determine a method for access. ATTN: Tim Kelley | Sailesh Gautam | Assigned |
| 7/22/2025 | 460375 | 0020 HOWE | This room is consistently warm, the air handlers do not seem to be moving any air. | Roger Suski | Assigned |
| 7/29/2025 | 461315 | 2624 HOWE | Lack of air flow and getting very warm | Laura Bestler | Open |
| 7/30/2025 | 461434 | 1140 HOWE | ** this door is in the northeast basement stairwell. It's the auditorium handicap entrance. *Safety issue.... looks like mold in carpet. This has been an on going problem for years. And work orders have been put in for leaks and flooding due to leaks and cracks in the concrete. They have fixed the flow of water outside. So it's hopefully stopped water from getting in. The carpet, ceiling tile, and vent covers needs to be replaced and weren't done in past work orders. Thanks in advance for replacing | Christi Johnson-Bascomb | Assigned |
| 8/1/2025 | 461726 | 0612 HOWE | Hire Midwest Automatic to repair existing 8" Watts 709 backflow preventer.: Basement Mechanical Room Fire Backflow: First check failed to hold it or above 1.0 psi. Recommend replacing spring and rubber gasket and retesting backflow. ATTN: Sailesh Gautam | Sailesh Gautam | Assigned |
| 8/11/2025 | 465331 | 1214 HOWE | Karcher extractor really needs a new vacuum head part that is pictured below as plastic stability tabs has one that is broken and the other is 1/2 fractured (duck taped in picture) that just started breaking turning on a rug I was trying to clean. | Nickolas Thielen | Materials-Awaiting |
| 8/20/2025 | 466503 | A HOWE | Repair vfd control wiring on AHU 6 attention Jeremy Forristall | Jeremy Forristall | Assigned |
| 8/21/2025 | 466726 | HOWE | The elevator on the west side of the building is stuck open on the 2nd floor. HOWE floor 2 | Andrew Albaugh | Assigned |
| 8/21/2025 | 466727 | 1280 HOWE | Brock- custodial equipment repair Pickup-Inside Dock door 2 blue carpet fans Plugged in and not turning on. Don't know if there is a reset button. Call me with questions. Gerdin fan was brought to me in July by accident and never turned on. Howe fan just stopped working today. We have air handlers out and it's been on non-stop for 2wks in atrium for air circulation. | Christi Johnson-Bascomb | Assigned |
| 8/21/2025 | 466728 | GENERAL HOWE | All 5 are- Can light fixtures Fixtures have been relamped Need new ballasts pls 1 in hallway intersection 0260 1 in hallway t-intersection 0040& 0200 2 above vending machine by restroom 0135 1 in hallway by restroom 0135 Thanks in advance for replacing | Christi Johnson-Bascomb | Assigned |
| 8/26/2025 | 467201 | 2270 HOWE | Can light over closet 2271 is not working properly. | Nickolas Thielen | Assigned |
| 8/29/2025 | 468155 | 1305 HOWE | Drinking fountain is leaking (located in the hallway next to 1305, the men's restroom). | Andrew Jordan | Materials-Awaiting |
| 9/10/2025 | 472260 | HOWE | Freight elevator west is not working/doors won't close Elevator # 10013548 Schumacher Invoice 90660891 Technician found a screw in the door track preventing the doors from closing. He removed the debris, checked overall operation and returned the car to service. .5 hour x $248.23 = $124.12 | Andrew Jordan | Open |
| 9/17/2025 | 474225 | 2624 HOWE | Room 2624 is very hot & humid | Laura Bestler | On Hold |
| 9/18/2025 | 474324 | 1635 HOWE | Brock Equipment fix Ivac spinner not spinning and is broke. Thanks in advance for fixing | Christi Johnson-Bascomb | Materials-Awaiting |
| 9/22/2025 | 474565 | 1305 HOWE | Cut an access hole in wall for Don and Scott | Donald Gregory | Assigned |
| 9/23/2025 | 474682 | HOWE | 2 leaks in the south mechanical room - Possibly the storm drains, please see John Breon for details. | John Breon | Assigned |
| 9/26/2025 | 474927 | 1140 HOWE | Please mark as critical: A significant number of lights are out in the Lee Liu Auditorium, making last night's midterm exam VERY challenging. We request this issue be addressed at least prior to next exam, mid-October. | Reagan Osgood |
Closed 11/3/2025
|
| 9/30/2025 | 475553 | 1140 HOWE | The auditorium lighting is nearly impossible to hold lectures in. This is also a space we host industrial partners in and I am going to start looking elsewhere because this is a disgrace. ¾ of the lights are out! Additional location information: Lee Liu Auditorium | Christine Nelson |
Closed 11/19/2025
|
| 10/1/2025 | 475660 | 0142 HOWE | Water is coming up through the floor tiles in the kitchen area. If you walk on the tiles, you can see moisture coming up between the cracks. This is also staining between the tiles in that kitchen area. Not sure if this is groundwater coming up, which we have not had rain in a long time, or something to do with a nearby closet that contains a sink that we do not have a key to. Once the water source has been identified and resolved, we need someone to look at whether those floor tiles need to be cleaned or replaced. It feels like the water may have loosened some of them. Room 142 is directly underneath the Howe Hall auditorium. | Glen Galvin | Assigned |
| 10/2/2025 | 478804 | 1365 HOWE | Expediency would be appreciated on this matter, so we can track down the odor before it goes away. There is a strange odor in room 1365. It does not seem to be coming from within the room. I suspect the odor is traveling through the HVAC into the room. The odor initially appeared this summer for a couple days and then went away. It came back a couple days ago and has persisted. We would like to track down its source before it goes away again. I work in a shop with a laser cutter, and to me it's the same smell as when we cut acrylic on the laser cutter. "Melted plastic". | Andrew Jordan | Assigned |
| 10/3/2025 | 478845 | 1131 HOWE | 1131 Howe Hall, men's bathroom, the hand towel dispenser on the right as you walk in no longer works. This is a high-traffic bathroom, so it would help to get that fixed | Stephen Gilbert |
Closed 11/10/2025
|
| 10/8/2025 | 479272 | 3200 HOWE | Repair electrical MCC vfd breaker tub | Jeremy Forristall | Assigned |
| 10/8/2025 | 479282 | 3200 HOWE | Repair vfd on supply2 on AHU 2. Attention Jeremy Forristall | Jeremy Forristall | Assigned |
| 10/17/2025 | 480057 | GENERAL HOWE | A piece of metal trim is falling off the ceiling by the skywalk on the east side. ATTN Gary Puls | Gary Puls |
Closed 10/31/2025
|
| 10/17/2025 | 480081 | 1638 HOWE | VAV-7-30 air flow sensor is not functioning properly. The measures SUP-FLOW in Metasys reades 0, which results in the Supply Damper opening up all the way. I have placed an Operator Override on the Supply Damper for now. Is it possible to clean or fix the Supply Airflow sensor so we can remove the override? This VAV is located along the south windows above the third cubicle from the east. The VAV-7-30 number is written on the unit in marker. Feel free to call me 4-6604 or stop by my office in 1620M for ladders or questions. | Glen Galvin | Assigned |
| 10/21/2025 | 480302 | 1140 HOWE | Request to test for mold - carpet. Contact Gail Koobs for more information. Work order associated to this is 461434 | Morgan Linley | Open |
| 10/24/2025 | 480693 | GENERAL HOWE | icing tunnel not controlling attn chris grosz | Christopher Grosz | Assigned |
| 10/29/2025 | 481270 | 1135 HOWE | This is the women's restroom. The middle left toilet seems to leak when you flush it. Water is collecting from the automatic flusher on the rim of the toilet. | Nicole Prestemon |
Closed 10/30/2025
|
| 10/30/2025 | 481347 | 0301 HOWE | toilet in the womens restroom is running | Laura Bestler |
Closed 10/31/2025
|
| 10/30/2025 | 481360 | 0301 HOWE | Toilet flushing non stop in women's restroom | Laura Guzman |
Closed 10/31/2025
|
| 10/30/2025 | 481379 | GENERAL HOWE | Please go to Rm 0635 and Nathan will go with you to the lab where there is a broken safety interlock on a 480 V disconnect. | Nathan Grambau |
Closed 11/17/2025
|
| 10/30/2025 | 481380 | GENERAL HOWE | Fixing padlock for roof access | John Breon |
Closed 10/30/2025
|
| 10/31/2025 | 481455 | 0609 HOWE | Replace UPS Batteries in Howe #03 Mx2 in room 609 | Steven Ristvedt |
Closed 11/3/2025
|
| 11/6/2025 | 485607 | 2644 HOWE | Could we please have someone shampoo the carpet in the space? If possible, we would prefer it to be done on a Friday to cause the least disturbance to staff. Work tags to attach: PG113293, DD03056 | Madeline Dickson | Assigned |
| 11/6/2025 | 485619 | 0305 HOWE | smaller toilet has severe crack on north side. | James Benson |
Closed 11/20/2025
|
| 11/6/2025 | 485643 | 1305 HOWE | The men's restroom door is locked from the inside | James Benson |
Closed 11/7/2025
|
| 11/7/2025 | 485671 | GENERAL HOWE | Assist professor with antenna work | John Breon | Open |
| 11/11/2025 | 485908 | 2332 HOWE | Single light bulb over cubicle station 2332-16 is not working properly | Nickolas Thielen |
Closed 11/14/2025
|
| 11/11/2025 | 485928 | 1131 HOWE | First urinal leaks every time it flushes | Malik Deloach |
Closed 11/13/2025
|
| 11/11/2025 | 485945 | 1200 HOWE | Water leak from ceiling and leaking on light fixture. Slow drip. Ceiling tile has been removed, and bucket placed to catch water. Have ladder by area. | James Benson |
Closed 11/14/2025
|
| 11/13/2025 | 486146 | 1305 HOWE | Door to men's restroom is about to fall off! | Nickolas Thielen |
Closed 11/18/2025
|
| 11/17/2025 | 486355 | 2233 HOWE | The handle to open the window has broken off, and he is unable to close the window. The window is currently stuck open. | Simone Servadio | Assigned |
| 11/18/2025 | 486411 | 2301 HOWE | 1st toilet from sinks is not flushing properly. | Nickolas Thielen |
Closed 11/21/2025
|
| 11/18/2025 | 486478 | 1350 HOWE | Howe Hall, located next to the back entrance, has a floor mat near 1380 that poses a tripping hazard. I've tried to straighten it out, but it's missing an edge. Jamie Niman's email address is jniman@iastate.edu new mat is ordered and will be put in place when its dropped off | Jamie Niman |
Closed 11/24/2025
|
| 11/18/2025 | 486482 | 1328C HOWE | 1328C Howe is currently hovering around 80°F, and we have an AV equipment rack along with 10 computers in this space. I would appreciate it if someone could check the thermostat for this room. Kyle D Solberg's email address is Kdsolb@iastate.edu | Kyle Solberg | Assigned |
| 11/19/2025 | 486517 | GENERAL HOWE | Howe Hall second floor, north side. All of the rooms are warmer than usual, and there is a lot of air pressure preventing doors from closing normally. Specific rooms checked include Howe 2232, 2234, and 2238. (I do not know if any other floors or sides of Howe Hall are experiencing this.) | Andrew Jordan | Assigned |
| 11/19/2025 | 486534 | 2232 HOWE | There is a burning smell in this room and there is a bulb burnt out in every fixture in the room. | Andrew Jordan | Assigned |
| 11/20/2025 | 486683 | HOWE | Attn: Burke Plain replace 6in valve That is leaking by | James Bohl |
Closed 11/24/2025
|
| 11/24/2025 | 486821 | 2271 HOWE | The rotating brush on the upright vacuum cleaner in the closet is broken. | Nickolas Thielen | Assigned |
| 11/24/2025 | 486841 | 2203 HOWE | 4th toilet stall from sink (closet one to outside door) is flushing excessively. | Nickolas Thielen |
Closed 11/24/2025
|
| 11/24/2025 | 486850 | 2305 HOWE | Left sink in men's restroom is dripping rapidly. | Nickolas Thielen |
Closed 11/24/2025
|
| 11/24/2025 | 486854 | 0612 HOWE | Heating water pump P1/ VFD1 is blowing the fuse on the step down transformer in the bottom of the drive. E025928 ATTN: Jeremy Forristall | Joseph Anastasi | Assigned |
| 11/24/2025 | 486855 | 0612 HOWE | Pump #3 reheat pump has very bad bearing noise and should be looked into and repaired. ATTN: Don Gregory | Joseph Anastasi | Materials-Awaiting |
Loading data, please wait...